Calculate the relatedness of a boy-child to a) his biological mother, b) stepfather, and c) full sister for i) a regular autosomal locus, and ii) for if he was a haplodiploid. Drawing the mother etc.’s genotypes and the gametes, indicate identity and non-identity with different colour lines, count ibd fractions and take a ratio to obtain the relatedness. Where the kin-value is important, calculate that too (20)


1
Expert's answer
2021-11-18T14:06:01-0500

= 1-(1/2×1/2)

= 1-(1/4)= 3/4